OpenStudy (anonymous):

Solve the following equation: 4x2 + 13x = -3

OpenStudy (zale101):

add 3 to both sides

OpenStudy (anonymous):

Okay, and what do i do afterwards?

OpenStudy (zale101):

show me your new equation after adding 3 to both sides

OpenStudy (anonymous):

4x2+13x+3=0

OpenStudy (zale101):

by using the standard form of quadratic, Ax^2+bx+c=0 which are replaced by 4x2+13x+3=0 use the quadratic formula to plug in A=4 B=13 C=3